OpenStudy (anonymous):

remove the bracket 4ab-[3{2(4a-b)+b(2-a)}]

OpenStudy (souvik):

3{2(4a-b)+b(2-a)} =3{2*4a-2*b+b*2-b*a} =3{8a-2b+2b-ba} =3{8a-ba} =3*8a-3*ba =24a-3ab

OpenStudy (anonymous):

4ab-[24a-3ab]

OpenStudy (souvik):

yes..

OpenStudy (anonymous):

is the possible answer would be 96a.a.b-12a.a.b.b? :D

OpenStudy (souvik):

4ab-[24a-3ab] =4ab-24a+3ab =7ab-24a

OpenStudy (anonymous):

sorry thanks!! :D

OpenStudy (souvik):

:)

OpenStudy (anonymous):

are you a math teacher?